OpenStudy (anonymous):

write the absolute value function as a piecewise function. y=I4x+1I+2x-3

terenzreignz (terenzreignz):

We need the 'turning' point of your absolute value function, that is, the value of x which would make the thing inside the bars || equal to zero. That said, when is 4x + 1 equal to zero?

OpenStudy (anonymous):

@terenzreignz when x=-1/4 ?

terenzreignz (terenzreignz):

Precisely :) So that's going to be the turning point of your function :D \[\Large f(x) = \left\{\begin{matrix}&\qquad x \le -\frac14\\\\\\\\\\&\qquad x> -\frac14\end{matrix}\right.\]

terenzreignz (terenzreignz):

Now, when x is less than -1/4 4x+1 is positive or negative?

OpenStudy (anonymous):

@terenzreignz negative?

terenzreignz (terenzreignz):

Yes. So, since it's negative, and in absolute value, it becomes positive when x is less than -1/4 Then we multiply it by -1 and it becomes -4x - 1 okay?

OpenStudy (anonymous):

ok that makes sense so far

terenzreignz (terenzreignz):

Right? Because |4x + 1| is basically the same as -4x - 1 whenever x is less than -1/4

terenzreignz (terenzreignz):

lol okay, so when x is less than or equal to -1/4, then the thing inside the absolute value gets multiplied by -1 (to turn it positive) \[\Large f(x) = \left\{\begin{matrix}\color{blue}{-4x-1}+2x -3&\qquad x \le -\frac14\\\\\\\\\\&\qquad x> -\frac14\end{matrix}\right.\] We will simplify later. Anyway, when x is greater than -1/4 then |4x + 1| is positive, yes?

terenzreignz (terenzreignz):

Rather, sorry (since |4x+1| is always positive anyway) when x is greater than -1/4 then 4x + 1 is positive, right?

OpenStudy (anonymous):

oh ok!!! so f(x)={ 6x-2 when x greater than or equal to -1/4 -2x-4 when x < -1/4

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Someone's eager :) LOL Correct :) \[\Large f(x) = \left\{\begin{matrix}-2x-4&\qquad x \le -\frac14\\\\\\\\\\6x-2&\qquad x> -\frac14\end{matrix}\right.\] Note that it doesn't matter if we switch the "or equal to" like this \[\Large f(x) = \left\{\begin{matrix}-2x-4&\qquad x \color{red}< -\frac14\\\\\\\\\\6x-2&\qquad x\color{red}{\ge}-\frac14\end{matrix}\right.\] Since both 'pieces' of the function have the same value when x = -1/4 Great job, by the way ^_^

OpenStudy (anonymous):

ok thank you so much! i finally understand this stuff now! :)

terenzreignz (terenzreignz):

Glad to hear it :) Signing off for now... ------------------------------ Terence out